The first ionisation energy of potassium is 100kcal mol-1.Calculate the lowest possible frequency of light that can ionise a potassium atom.

Grade:12

1 Answers

Mahima Kanawat
6 years ago
Dear student .....To ionise a K atom we require minimum 100 kcal ( 100×10³×4.2 J) energy.......As we know E=hv.....42×10⁴=6.6×10³⁴×v.....v=6.3×10^-30Hz......With Regards .....Mahima..(Askiitians forum expert )